OpenStudy (ashley1nonly):

) A reaction is found to have an activation energy of 38.0 kJ/mol. If the rate constant for this reaction is 1.60 × 102 M-1s-1 at 249 K, what is the rate constant at 436 K? A) 2.38 × 105 M-1s-1 B) 1.26 × 103 M-1s-1 C) 7.94 × 104 M-1s-1 D) 4.20 × 105 M-1s-1 E) 3.80 × 104 M-1s-1 Answer: D

OpenStudy (ashley1nonly):

we will use lnk2/k1=( Ea/R)(1/t1-1/t2)

OpenStudy (anonymous):

kk 1 sec

OpenStudy (ashley1nonly):

ok

OpenStudy (ashley1nonly):

Im working in a group preparing for my chemistry exam and my group wants to just praise to you for helping us understand these problems. we have a test on Monday and you made everything so simple and we will pass now lol

OpenStudy (anonymous):

Im happy to hear that I could help, and you will all do excellent

OpenStudy (anonymous):

kk I got this one

OpenStudy (ashley1nonly):

lets do it

OpenStudy (anonymous):

1. convert 38.0kj to joules so 38*10^3Joule R=constant 8.3145 2.equation 1.60^2*e^(38Jmole*8.3145)*(249-436) 3. 1.60^2e^(7.87233)=4.20*105

OpenStudy (ashley1nonly):

what variable are we solving for so i can mentally see where you are going with the numbers

OpenStudy (anonymous):

kk 1 sec

OpenStudy (anonymous):

we are solving for k2

OpenStudy (ashley1nonly):

working it out to see if i can get the answer

OpenStudy (anonymous):

1.) ln(k2/k1)=(Ea/R)(t1-t2) 2.) ln(1.60*10^2/k1)=(38*10^3)(249-436) 3.) ln(1.60*10^2/k1)(=(65.4544) 4 k1(1.60*10^2) =65.4544/ln 5.k1=1.60*10^2*e^65.4544= 4.20*105 Remember when dividing by ln, make it e^

OpenStudy (ashley1nonly):

so we were solving for k1 or k2

OpenStudy (anonymous):

basically just use algebra rules

OpenStudy (ashley1nonly):

okay that makes sense now

OpenStudy (ashley1nonly):

okay we have one more which is similar to this one

OpenStudy (anonymous):

is it making sense, sorry I am a slow typer

OpenStudy (ashley1nonly):

yes the problem we had was getting rid of the ln. we had no clue

OpenStudy (ashley1nonly):

also it is easier to fully solve the right side of the equation and just have one number instead of all the equations

OpenStudy (ashley1nonly):

then once we do that we can take the e to get rid of the ln and the multiply to get rid of the k1 which leaves us with k2

OpenStudy (anonymous):

Basically for this problem, divide the right side with ln (ea*ti-t2) make it say e^value

OpenStudy (ashley1nonly):

yes

OpenStudy (ashley1nonly):

im going to post the last question that we have for you and that will complete us lol

OpenStudy (anonymous):

yes k2 is what you are finding k1=1.60*10^2

OpenStudy (anonymous):

kk sounds good, did this problem make sense

OpenStudy (ashley1nonly):

at first it didnt because i didnt know what we were looking for and i got lost in all of the different variables but the way you broke it down shows all of the steps i needed.

OpenStudy (anonymous):

kk awseome!